Calculate the magnitude of the force of repulsion between two equal charges of 2.0V separated by a distance of 1m.

Take (1/4"\lambda\epsilon\Omicron"=9x109)

1
Expert's answer
2022-03-28T11:02:39-0400

The Coulomb's says

"F=k\frac{q_1q_2}{r^2}=9*10^9*\frac{2.0*2.0}{1^2}=36*10^9\:\rm N"